Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Roma Mauro at Fri Feb 22 09:06:07 MET 2002) Minor failures
All scan is recorded. From start time to 01:00UT wind and rain.

Torun (K.Borkowski at Fri Feb 22 09:10:07 MET 2002) Success
No known problems (the wavelength entry in the feedback page do not apply).

Jodrell_Bank (Alastair Gunn at Fri Feb 22 10:12:39 MET 2002) Minor failures
Telescope parked due to high winds at 07:04 UT on 22/02/02 until end of experiment. No other known problems.

Westerbork (Tony Foley at Fri Feb 22 16:13:46 MET 2002) Minor failures
Lost approximately 04:30 to 05:00 and 06:00 to 08:00 UT with online control software problems (we think).

Effelsberg (A. Kraus at Mon Feb 25 08:38:31 MET 2002) Minor failures
We had to stop between UT1700-1715 (computer problems), UT1801-1807 and UT1828-1832 (azimuth drive problems), and from UT0319-end (strong winds and snow).

\(Other\) (John Conway at Tue Feb 26 17:12:37 MET 2002) Minor failures
Telescope stowed after 19:17 UT due to high winds so lost second half of expt.

Onsala (John Conway at Tue Feb 26 17:13:26 MET 2002) Minor failures
Above comment was for Onsala

Westerbork (Tony Foley at Fri Mar 1 11:50:53 MET 2002) Minor failures
using array of 12 telescopes - missing RTC and RTD Large sections from 22 03 to 23;05 out of hour angle range.